# Benchmarking Language Model Creativity: A Case Study on Code Generation

### v1

@@ 6 Conclusion
We propose two key protocols for evaluating language model creativity and introduce the Denial Prompting framework and NeoGauge metric to provide a one-stop creativity evaluation that follows the protocols. Our NeoGauge stands out for its comprehensive evaluation of both convergent and divergent creativity, aligning with concepts from human creativity studies. To facilitate future research, we release our NeoCoder dataset and shed light on the limitations of current reasoning strategies in improving LLM creativity.
@@ Acknowledgements
Appendix A Additional Details of Experimental Setup
Appendix B Additional Details of Experimental Results
Appendix C Prompts for Denial Prompting and Benchmarking
Appendix A Experiment Setup
A.1 Human Creativity Evaluation
@@
Given total $n$ problems, where each problem has $m$ human solutions, we compute the average ratio of new techniques used by a single human solution $\hat{y}_{j}^{i}$ ($j^{\text{th}}$ human solution for $i^{\text{th}}$ problem) that the remaining human solutions $\{\hat{y}_{k}^{i}\mid k\neq j,k\ =1,2,\cdots,m\}$ have never used. The only difference to Eq.3 is that because it is not practical to ask human coding experts to perform Denial Prompting on hundreds of programming problems, human divergent creativity in Eq.6 is thus state insensitive and is equivalent to divergent(human, $t=0$).
Appendix B Experiment Results
@@
We plot the stacked convergent and divergent creativity evaluation results in Fig.8. Among all models, GPT-4 generally exhibits the best performance on both convergent and divergent creative thinking across all states, followed by Claude-3 and Llama3-70B. It is noticeable that Llama3-70B even outperforms GPT-4 on convergent creative thinking when $t=0$ (convergent(GPT-4, 00) $=16.16$ < convergent(Llama3-70B, 00) $=19.19$). We hypothesize that the latest Llama3 models are pre-trained on Codeforces problems and human solutions, so they have superior performance when there is no external constraint $t=0$. However, as $t$ increases, its convergent performance drops drastically. Moreover, divergent creative thinking never goes to 0 across all states and is sometimes even equally distributed on those less small models (e.g., CodeGemma-7B and Mistral-7B). Together with independent findings from Xu et al. [57], this observation indicates that LLMs with insufficient reasoning capabilities tend to make up new solutions regardless of the quality when facing unusual problems. Which, in turn, demonstrates the importance of the claim we made in section 1 that creative thinking involves not merely the generation of many diverse alternatives but also the verification of new valid alternatives.
Appendix C Prompts for Denial Prompting and Benchmarking
Technique Dection Prompt: You are a code reviewer. Detect all the programming techniques from the input and return a list of programming techniques. Only select the techniques from this list: [’if statement’, ’for loop’, ’while loop’, ’break statement’, ’continue statement’, ’pass statement’, ’match statement’, ’recursion’, ’stack’, ’queue’, ’tuple’, ’set’, ’dictionary’, ’linked list’, ’tree’, ’graph’, ’graph traversal’, ’two pointers’, ’sliding window’, ’matrix operation’, ’hashmap’, ’depth first search’, ’width first search’, ’back tracking’, ’dived & conquer’, ’Kadanes algorithm’, ’binary search’, ’heap’, ’dynamic programming’, ’greedy algorithm’, ’misc’, ’minimax’, ’topological sort’, ’sorting’, ’graph traversal’]
Your output should look like this:
- technique 1
- technique 2
- technique 3
State Constraint Problem Statement
[…]

### v2

@@ 6 Conclusion
We propose protocols for evaluating language model creativity in problem-solving and introduce the Denial Prompting framework and NeoGauge metric to provide a comprehensive creativity evaluation, measuring both convergent and divergent creativity, inspired by extensive research on human creativity. To facilitate future research, we release our NeoCoder dataset and shed light on the limitations of current reasoning strategies in improving LLM creativity.
@@ Acknowledgements
Appendix A Why Choose Codeforces for Creativity Evaluation
Appendix B Additional Details of Experimental Setup
Appendix C Additional Details of Experimental Results
Appendix D Prompts for Denial Prompting and Benchmarking
Appendix A Why Choose Codeforces for Creativity Evaluation?
In this study, we use competitive programming problems sourced from Codeforces for creativity evaluation. We provide our task choice motivation by answering the following three interrelated questions.
Why choose competitive programming problems?
The general purpose of this paper is to benchmark the LLM’s creativity performance in dealing with unconventional and challenging problems. Understandably, these problems usually do not have ground-truth answers (e.g., how to make coffee without a coffee maker). In such cases, we typically either evaluate the generated solution through human evaluation, similar to the approach taken by Tian et al. (2024), or through automated machine evaluation (ours). Real-world problems (Tian et al., 2024) naturally need human annotation. Collecting human annotations for measuring machine creativity is particularly challenging since the space is typically vast (because of the nature of creativity). Conversely, coding becomes an ideal source for problems that can be its functional correctness (as opposed to the choice of syntax) evaluated automatically with a minimal cost—based on whether they pass the test cases. Thus, we first chose coding problems to examine LM’s creativity, as they provide an open-ended environment that could stimulate a model’s creativity performance while making evaluation easy and cost-effective.
Low performance or low creativity?
The low pass rate and constraint following ratio in Table 3 may raise a new question as to whether there are no reasonable solutions at all or no requisite creativity in finding solutions. Experimental evidence, however, suggests that LM simply lacks creativity. According to Figure 5, the huge gap between human and LLMs convergent creativity prior to State 3 (0-3 constraints) indicates there are valid human solutions for each problem, but the LLMs seem to be lacking creativity in finding it. Additionally, according to Figure 6, with suitable reasoning strategies, LLM still has room for improvement in both convergent and divergent creativity. Even though humans’ convergent scores are nearing zero (Figure 5) at a large state (>3 hard constraints), the problems might not be fully infeasible.
Why not evaluate creativity based on problems but solutions?
A motivational example for this question is that a creative student can always come up with innovative and insightful questions. However, in this work, we adopt a different standpoint on creativity used by many psychological and cognitive studies (discussed in section 2), which emphasizes problem-solving abilities. We evaluate a student to be creative if he/she can leverage all available tools and come up with novel solutions for challenging problems. Similarly, we study LLM creativity based on solutions they generated for challenging programming problems.
Appendix B Experiment Setup
B.1 Human Creativity Evaluation
@@
Given total $n$ problems, where each problem has $m$ human solutions, we compute the average ratio of new techniques used by a single human solution $\hat{y}_{j}^{i}$ ($j^{\text{th}}$ human solution for $i^{\text{th}}$ problem) that the remaining human solutions $\{\hat{y}_{k}^{i}\mid k\neq j,k\ =1,2,\cdots,m\}$ have never used. This is because collecting a human DP dataset would be quite costly and restrictive. We instead use a diverse collection of solutions from various human programmers as a proxy. Eq.6 is equivalent to divergent(human, $t=0$), representing the lowest level of human divergent creativity.
Appendix C Experiment Results
@@
We plot the stacked convergent and divergent creativity evaluation results in Figure 7. Among all models, GPT-4 generally exhibits the best performance on both convergent and divergent creative thinking across all states, followed by Claude-3 and Llama3-70B. It is noticeable that Llama3-70B even outperforms GPT-4 on convergent creative thinking when $t=0$ (convergent(GPT-4, 00) $=16.16$ < convergent(Llama3-70B, 00) $=19.19$). We hypothesize that the latest Llama3 models are pre-trained on Codeforces problems and human solutions, so they have superior performance when there is no external constraint $t=0$. However, as $t$ increases, its convergent performance drops drastically. Moreover, divergent creative thinking never goes to 0 across all states and is sometimes even equally distributed on those less small models (e.g., CodeGemma-7B and Mistral-7B). Together with independent findings from Xu et al. (2024b), this observation indicates that LLMs with insufficient reasoning capabilities tend to make up new solutions regardless of the quality when facing unusual problems. Which, in turn, demonstrates the importance of the claim we made in section 1 that creative thinking involves not merely the generation of many diverse alternatives but also the verification of new valid alternatives.
Appendix D Prompts for Denial Prompting and Benchmarking
Technique Dection Prompt: You are a code reviewer. Detect all the programming techniques from the input and return a list of programming techniques. Only select the techniques from this list: [’if statement’, ’for loop’, ’while loop’, ’break statement’, ’continue statement’, ’pass statement’, ’match statement’, ’recursion’, ’stack’, ’queue’, ’tuple’, ’set’, ’dictionary’, ’linked list’, ’tree’, ’graph’, ’two pointers’, ’sliding window’, ’matrix operation’, ’hashmap’, ’depth first search’, ’width first search’, ’back tracking’, ’divide & conquer’, ’Kadanes algorithm’, ’binary search’, ’heap’, ’dynamic programming’, ’greedy algorithm’, ’misc’, ’minimax’, ’topological sort’, ’sorting’, ’graph traversal’]
Your output should look like this:
- technique 1
- technique 2
- technique 3
State Constraint Problem Statement
[…]